Grade 11Mechanics

A partical move 20m in 2s towards east then it moves 4m in 3s towards south then find average speed

Niharika Malviya
6 Years agoGrade 11

2 Answers

Arun
6 Years ago
Dear student 
 
Average speed = total distance / total time taken
 
 = 24/5 = 4.8 m/sec
 
Hope it helps
 
Regards
Vikas TU
6 Years ago
Dear student 
Speed is a scalar quantity , So we can not take direction in case of scalar quantity . 
If it would be velocity then we will solve it by using magbnitude and directon. 
Total distance travelled = 20 + 4 = 24 meter 
Total time = 5 sec 
Avg speed = Total distance / total time 
Avg speed = 24 / 5 = 4.8 m/sec 
Hope this clears 
Good Luck